OpenStudy (hellokitty17):

Write x2 - 2x - 3 = 0 in the form (x - a)2 = b, where a and b are integers. (x - 4)2 = 3 (x - 3)2 = 2 (x - 2)2 = 1 (x - 1)2 = 4

oregonduck (oregonduck):

We have: x^2 - 2x - 3 = 0 To complete the square, we take the middle number, halve it and then square it. (-2 / 2)^2 = 1 Now we have the answer, '1', we add this and take it to the end of the equation: x^2 - 2x + 1 - 1 - 3 = 0 This then allows us to simplify the first part as a perfect square: x^2 - 2x + 1 - 1 - 3 = 0 (x - 1)^2 - 1 - 3 = 0 (x - 1)^2 = 4 So the answer is: (x - 1)^2 = 4 ANSWER IS D